A driver at Dewaan Motors is testing a new model car with a speedometer calibrated to read m/s

rather than km/h. The following series of speedometer readings was obtained during a test run

along a long, straight road:

Time (s)

Speed (m/s)

0 1 3 5 7 9

0 0 4 7 10 13

15

17

19

11 13

17 22

23

26

29

(a) Compute the average acceleration during each 2-s interval. Is the acceleration constant? Is it

constant during any part of the test run?

(b) Make a (v-t) graph of the data, using scales of 1cm = 1s horizontally and 1cm = 2m/s

vertically. Draw a smooth curve through the plotted points. By measuring the slope of your

curve, find the instantaneous acceleration at t = 6 s, 18 s, and 25 s.


1
Expert's answer
2021-01-27T14:28:45-0500

a)




Acceleration is not a constant.

It is constant from 3 s to 9 s, and from 17 s to 19 s.

b)


"a(6)=a(18)=a(25)=1.5\frac{m}{s^2}"